OpenStudy (anonymous):

find all numbers for which the rational expression is not defined: b -5/5b-10

OpenStudy (mathmath333):

is this the original equation \(\Large \dfrac{b -5}{5b-10}\)

OpenStudy (anonymous):

yes that is the original one.

OpenStudy (anonymous):

honestly i want to know how to solve for this.

OpenStudy (mathmath333):

well the equation we want is \(\Large \dfrac{b -5}{5b-10}=\infty\)

OpenStudy (mathmath333):

that is \(\Large \dfrac{b -5}{5b-10}=\dfrac{1}{0}\)

OpenStudy (mathmath333):

for that we need to make denominator \(\Large 5b-10=0\)

OpenStudy (mathmath333):

hence for \(\Large b=2\) the original equation is indefined

OpenStudy (anonymous):

ok. that is awesome! thank you so much for explaining how to do this.

OpenStudy (mathmath333):

welcome